The demand and supply for a good are described by the following functions: Q(p) = 3000 – 10p and Q(p) = 2000 + 10p. If Government impose a tax of RON 20 for each unit of product sold, the tax burden will be shared:


1
Expert's answer
2021-01-27T09:13:59-0500

1) Let's find equilibrium price and quantity before tax imposing: 3000-10p=2000+10p; p= 50; q=2500

2) Let's find equilibrium price and quantity after tax imposing: 3000-10p=2000+ 10(p-20); p= 60, q=2400

3) The government revenue is 2400×20=48000. It consists of consumer's and producer's share. The consumer's burden is the decrease in consumers surplus. The consumer's surplus before tax is 1/2×(300-50)×2500=312500 (where 300 is the maximum price at which q=0). The consumer's surplus after tax is 1/2×(300-60)×2400=288000. So, consumers' burden is 312500-288000=24500. The rest is producers' burden:48000-24500=23500.